OpenStudy (anonymous):

2/3 = 8/x A. 2, x B. 2, 3 C. 3, 8 D. 8, x

OpenStudy (anonymous):

2/3=8/x 2*x=3*8 x=3*8/2 x=3*4 x=12

OpenStudy (anonymous):

your working is ok but what is the question? the choices don't match